# MA4DIV: Multi-Agent Reinforcement Learning for Search Result Diversification

### v2

@@ Abstract.
The objective of search result diversification (SRD) is to ensure that selected documents cover as many different subtopics as possible. Existing methods primarily utilize a paradigm of ”greedy selection”, i.e., selecting one document with the highest diversity score at a time. These approaches tend to be inefficient and are easily trapped in a suboptimal state. In addition, some other methods aim to approximately optimize the diversity metric, such as $\alpha$-NDCG, but the results still remain suboptimal. To address these challenges, we introduce Multi-Agent reinforcement learning (MARL) for search result DIVersity, which called MA4DIV. In this approach, each document is an agent and the search result diversification is modeled as a cooperative task among multiple agents. This approach allows for directly optimizing the diversity metrics, such as $\alpha$-NDCG, while achieving high training efficiency. We conducted preliminary experiments on public TREC datasets to demonstrate the effectiveness and potential of MA4DIV. Considering the limited number of queries in public TREC datasets, we construct a large-scale dataset from industry sources and show that MA4DIV achieves substantial improvements in both effectiveness and efficiency than existing baselines on a industrial scale dataset.
@@ 1. Introduction
How to provide diverse search results to users is an important yet complicated problem in information retrieval. The search result diversification (SRD) (Agrawal et al., 2009) means that for a user’s search query, we need to provide search results that cover a wide range of subtopics to meet the requirements of users who may have multiple interpretations or intentions. Many diversity-aware metrics, such as $\alpha$-NDCG (Clarke et al., 2008), ERR-IA (Chapelle et al., 2011), S-recall (Zhai et al., 2015), have been proposed to evaluate the effectiveness of search result diversification. Since the diversity of a given document is affected by the documents that precede it, optimizing some of the diversity metrics is an NP-Hard problem (Carterette, 2011). Therefore, the diversity of search results has become an important and challenging research topic.
@@
The typical search result diversification method includes heuristic methods and learning-based methods. These methods use ”greedy selection” to sequentially select documents to build a diversified ranking list in a step-by-step manner. At each step, the ranking model selects the best document for the current ranking position according to the additional utility it can bring to the whole ranking list. Different methods use different criteria in the greedy selection. Maximal marginal relevance (MMR) (Carbonell and Goldstein, 1998) algorithm proposes maximal marginal, which is defined as the sum of query-document dependencies, and takes the maximum document distance as the utility. xQuAD (Santos et al., 2010) is another widely used diverse ranking model which defines the utility to explicitly account for relationship between documents retrieved for the original query and the possible sub-queries. In recent years, many methods based on machine learning (Zhu et al., 2014; Xu et al., 2017; Li et al., 2009; Mihalkova and Mooney, 2009; Qin et al., 2023) have been proposed to solve the search result diversification task. For example, the relational learning to rank (R-LTR) (Zhu et al., 2014) and its variants utilize a linear combination of correlation features and novelty features.
@@
All above methods focus on estimating the utility of candidate documents directly based on user queries and previously selected documents, but fail to explicitly model and fully utilized the utility that the user has gain in examining previous documents. Xia et al. (2017) propose the MDP-DIV framework to model the search result diversification ranking as Markov Decision Process (MDP) (Puterman, 2014), and uses a reinforcement learning algorithm (Williams, 1992) to optimize the ranking model. While MDP-DIV still use a ”greedy selection” to select next candidate result, it can easily lead to sub-optimal ranking list. $\mathrm{M^{2}Div}$ (Feng et al., 2018) introduces Monte Carlo Tree Search (MCTS) (Coulom, 2006) for high-quality exploration. This method helps to alleviate the problem of being stuck in a sub-optimal solution at a cost of increasing training and inference time.
@@
Besides these greedy selection methods that sequentially score and select each candidate results, some recent studies DALETOR (Yan et al., 2021) and MO4SRD (Yu, 2022) propose to simultaneously score all candidate documents and obtain the whole ranking list by sorting these scores in a descending order. This paradigm can greatly improve the inference efficiency as the documents can be scored in parallel. In addition, DALETOR and MO4SRD propose a differentiable proxy for the non-differentiable diversity metric $\alpha$-NDCG (Clarke et al., 2008), which enables a gradient-based optimization for the diversity metric. However, since the optimization goal is an approximation to the real evaluation metric, the solution may still be sub-optimal.
@@
To summarize, although the problem of search result diversification has been extensively investigated, existing methods still have some limitations in terms of the effectiveness and efficiency. From the perspective of effectiveness, the heuristic-based methods (e.g. (Carbonell and Goldstein, 1998; Santos et al., 2010)) and the approximation-based methods (e.g.(Yan et al., 2021; Yu, 2022)) may be sub-optimal in optimizing the end diversity metric, such as $\alpha$-NDCG and ERR-IA. From the perspective of efficiency, those RL-based approaches face a intrinsic challenge of exploring a huge space of all possible rankings, which inevitably leads to sub-optimal solution (Xia et al., 2017) or high time-complexity in both training and inference (Feng et al., 2018).
@@
In this paper, in order to alleviate the problems of existing methods mentioned above, we propose to formalize the diverse ranking as a multi-agent cooperation process. Since Multi-Agent reinforcement learning (MARL) algorithm is used to optimize the DIVersity ranking model, we call this new method MA4DIV. Specifically, by treating each document as an independent agent within a Decentralized Partially Observable Markov Decision Process (Dec-POMDP) setting, we simulate a fully cooperative multi-agent task. In this setup, each agent (document) makes action selections based on observations that include the features of both the query and the documents, aiming to maximize a shared reward function that is directly related to the diversity evaluation metrics of search results. Considering permutation equivariant (Pang et al., 2020) property of the ranking task, our MA4DIV framework leverages the Multi-Head Self-Attention (MHSA) for capturing the intricate relationships between documents. Furthermore, MA4DIV employs value decomposition (Rashid et al., 2018) to optimize global diversity directly by the structures of mixing network and hypernetworks during training.
@@
(1) Compared to ”greedy selection” paradigm, MA4DIV simultaneously predicts the ranking scores of all documents, which can improve the efficiency of ranking process and alleviate the local optimum problem.
(2) The ranking scores of all documents and a ranked documents list with diversity can be obtained in one time step. Therefore, MA4DIV has higher exploration efficiency than single-agent RL algorithms, such as MDP-DIV and $\mathrm{M^{2}Div}$, which must go through an entire episode to explore a diversity ranking list.
(3) The training of the MA4DIV model does not require any approximations of different diversity metrics as they can be directly optimized as reward in the multi-agent reinforcement learning process.
In order to demonstrate the effectiveness and efficiency of MA4DIV, we conducted experiments on TREC benchmark datasets and Baidu dataset. The experimental results on TREC datasets showed that, MA4DIV achieved the state-of-the-art performance on some evaluation metrics, and training time was significantly shorter than baselines. Considering the small number of queries in TREC dataset (only 198 valid queries), we built a larger diversity dataset based on the real Baidu search engine data, which is called DU-DIV. On the DU-DIV dataset, MA4DIV achieved the state-of-the-art performance on all evaluation metrics, and demonstrated a substantial improvement in exploration efficiency and training efficiency.
2. RELATED WORK
@@ 2.1. Search Result Diversification
The construction of a diverse ranking list in search is commonly formalized as a sequential document selection process or a process of greedy sequential decision-making. The aim is to design effective criteria for estimating the utility a document can bring, which is believed to depend not only on the document itself but also on the preceding documents. Various models have been proposed to guide this selection.
Carbonell and Goldstein introduced the maximal marginal relevance criterion (Carbonell and Goldstein, 1998), which uses a linear combination of query-document relevance and document novelty to determine the document to be selected. This process is iterative, with the marginal relevance score updated as the number of documents in the result set increases. An extension of this concept, the probabilistic latent MMR model (Guo and Sanner, 2010), was proposed by Guo and Sanner.
@@
Approaches such as xQuAD (Santos et al., 2010) directly model different aspects of a query, estimating utility based on the relevance of retrieved documents to identified sub-queries or aspects. Hu et al. (Hu et al., 2015) proposed a utility function that leverages hierarchical intents of queries, aiming to maximize diversity in the hierarchical structure. Other researchers, like He et al. (He et al., 2012), have proposed combining implicit and explicit topic representations to create better diverse rankings, while Gollapudi and Sharma (Gollapudi and Sharma, 2009) proposed an axiomatic approach to result diversification.
@@
Machine learning techniques have also been applied to construct diverse ranking models, often adhering to the sequential document selection or greedy sequential decision-making framework. These techniques aim to automatically learn the utility function based on training queries. Several researchers have defined utility as a linear combination of handcrafted relevance and novelty features (Zhu et al., 2014; Xia et al., 2015; Xu et al., 2017). In (Xia et al., 2016), novelty can be modeled using deep learning models such as neural tensor networks. Radlinski et al. (2008) propose to learn a diverse ranking of documents directly based on users’ clicking behavior.
@@
Reinforcement learning algorithms also are proposed to improve the effectiveness of ”greedy selection” paradigm in search result diversification. MDP-DIV (Xia et al., 2017) consider the diverse ranking process as a Markov Decision Process (MDP). In MDP-DIV, agent select one document in one time step, so a ranked list can be obtained after an episode. And the episode reward can be any metrics, such as $\alpha$-DCG. While Feng et al. (2018) think that ”greedy selection” paradigm is easily lead to local optimum, so $\mathrm{M^{2}Div}$ is proposed to introduce Monte Carlo Tree Search (MCTS) into MDP ranking process. In this way, $\mathrm{M^{2}Div}$ does alleviate the problem that tend to fall into local optimum by conducting high-quality exploration, but also brings expensive training and inference costs.
@@
There are other methods which approximate the diversity evaluation metrics to be differentiable and optimize the approximated metrics directly through the loss function. Based on score-and-sort paradigm, DALETOR (Yan et al., 2021) and MO4SRD (Yu, 2022) are proposed to derive differentiable diversification-aware losses which are approximation of different diversity metrics, such as $\alpha$-NDCG.
@@ 2.2. Reinforcement Learning for IR
Xu et al. (Wei et al., 2017) formulate a ranking process as an MDP and train the ranking model with a policy gradient algorithm of REINFORCE (Williams, 1992). Singh et al. (Singh and Joachims, 2019) propose PG Rank algorithm to enhance the fairness of ranking process with reinforcement learning methods. PPG (Xu et al., 2020) uses pairwise comparisons of two sampled document list with in a same query, which makes an unbiased and low variance policy gradient estimations. Yao et al. (Yao et al., 2020) model the interactions between search engine and users with a hierarchical Markov Decision Process, which improves the personalized search performance. RLIRank (Zhou and Agichtein, 2020) implement a learning to rank (LTR) model for each iteration of the dynamic search process. Zou et al. (2019) formulate the ranking process as a multi-agent Markov Decision Process, where the interactions among documents are considered in the ranking process. However, Zou et al. (2019) conduct the single-agent reinforcement learning algorithm REINFORCE to optimize the ranking model.
@@ 3. Background
3.1. Dec-POMDP
In this work, we consider the process of diversity ranking as a fully cooperative multi-agent task which is usually modeled as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP) (Oliehoek and Amato, 2016). Dec-POMDP is formally defined as a tuple $G=\left\langle\mathcal{S},\mathcal{A},\mathcal{P},r,\mathcal{Z},\mathcal{O},n,\gamma\right\rangle$. $s\in\mathcal{S}$ is the global state of the environment. Each agent $i\in\mathcal{G}\equiv\{1,\dots,n\}$ chooses an action $a^{i}\in A$ which forms the joint action $\textbf{a}\in\mathcal{A}\equiv A^{n}$. The state transition function is modeled as $\mathcal{P}(s^{\prime}|s,\textbf{a}):\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$. The reward function which is modeled as $r(s,\textbf{a}):\mathcal{S}\times\mathcal{A}$ is shared by all agents and the discount factor is $\gamma\in[0,1)$. It follows partially observable settings, where agents do not have access to the global state. Instead, it samples observations $z\in\mathcal{Z}$ according to observation function $\mathcal{O}(s,i):\mathcal{S}\times\mathcal{A}\to\mathcal{Z}$. Each agent has an action-observation history trajectory $\tau^{i}\in T\equiv(\mathcal{Z}\times\mathcal{A})^{*}$, on which it conditions a stochastic policy $\pi^{i}(u^{i}|\tau^{i}):T\times\mathcal{A}\to[0,1]$. In our algorithm, the joint policy $\pi$ is based on action-value function $Q^{\pi}_{tot}(s_{t},\textbf{a}_{t})=\mathbb{E}_{s_{t+1}:\infty,\textbf{a}_{t+1}:\infty}[\sum_{k=0}^{\infty}\gamma^{k}r_{t+k}|s_{t},\textbf{a}_{t}]$. The final goal is to get the optimal action-value function $Q^{*}$.
3.2. Diversity Evaluation Metrics
3.2.1. $\alpha\textrm{-}NDCG$
3.2.2. $ERR\textrm{-}IA$
The $ERR\textrm{-}IA$ is defined in Equation (3).
3.2.3. $S\textrm{-}recall$
3.3. Test Collection for Diversified Search

### v3

@@ Abstract.
Search result diversification (SRD), which aims to ensure that documents in a ranking list cover a broad range of subtopics, is a significant and widely studied problem in Information Retrieval and Web Search. Existing methods primarily utilize a paradigm of ”greedy selection”, i.e., selecting one document with the highest diversity score at a time or optimize an approximation of the objective function. These approaches tend to be inefficient and are easily trapped in a suboptimal state. To address these challenges, we introduce Multi-Agent reinforcement learning (MARL) for search result DIVersity, which called MA4DIV  [^1] The code of MA4DIV can be seen on https://github.com/chenyiqun/MA4DIV.. In this approach, each document is an agent and the search result diversification is modeled as a cooperative task among multiple agents. By modeling the SRD ranking problem as a cooperative MARL problem, this approach allows for directly optimizing the diversity metrics, such as $\alpha$-NDCG, while achieving high training efficiency. We conducted experiments on public TREC datasets and a larger scale dataset in the industrial setting. The experiemnts show that MA4DIV achieves substantial improvements in both effectiveness and efficiency than existing baselines, especially on the industrial dataset.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3696410.3714862isbn: 979-8-4007-1274-6/25/04ccs: Information systems Information retrieval diversity
@@ 1. Introduction
How to provide diverse search results to users is an important yet complicated problem in information retrieval. The search result diversification (SRD) (Agrawal et al., 2009) means that for a user’s search query, we need to provide search results that cover a wide range of subtopics to meet the requirements of different users who may have multiple interpretations or intentions. Many diversity-aware metrics, such as $\alpha$-NDCG (Clarke et al., 2008), ERR-IA (Chapelle et al., 2011), S-recall (Zhai et al., 2015), have been proposed to evaluate the effectiveness of search result diversification. Since the diversity of a given document is affected by the documents that precede it, optimizing some of the diversity metrics is NP-Hard (Carterette, 2011). Therefore, search result diversification has become an important and challenging research topic.
Existing works on SRD can primarily be categorized into three main approaches:
@@
Greedy Selection Approaches  The typical search result diversification methods use ”greedy selection” to sequentially select documents to build a diversified ranking list in a step-by-step manner. At each step, the ranking model selects the best document for the current ranking position according to the additional utility it can bring to the whole ranking list. Different methods use different criteria in the greedy selection. Maximal marginal relevance (MMR) (Carbonell and Goldstein, 1998) algorithm proposes maximal marginal, and takes the maximum document distance as the utility. xQuAD (Santos et al., 2010) is another widely used diverse ranking model which defines the utility to explicitly account for relationship between documents and the possible sub-queries. In recent years, many methods based on machine learning (Zhu et al., 2014; Xu et al., 2017; Li et al., 2009; Mihalkova and Mooney, 2009; Qin et al., 2023) have been proposed to solve the search result diversification task.
Single-Agent Reinforcement Learning Approaches
@@
While the greedy selection approaches only consider the myopic utility brought by each document, RL approaches can better models the current state (i.e., the documents examined by the user) and future expected reward. Xia et al. (Xia et al., 2017) propose the MDP-DIV framework to model the search result diversification ranking as Markov Decision Process (MDP) (Puterman, 2014), and uses a reinforcement learning (RL) algorithm (Williams, 1992) to optimize the ranking model. While MDP-DIV still use a ”greedy selection” to select next candidate result, it can easily lead to sub-optimal ranking list. $\mathrm{M^{2}Div}$ (Feng et al., 2018) introduces Monte Carlo Tree Search (MCTS) (Coulom, 2006) for high-quality exploration. This method helps to alleviate the problem of being stuck in a sub-optimal solution at a cost of increasing training and inference time.
@@
Simultaneous Scoring and Ranking Approaches  Some recent studies propose to simultaneously score all candidate documents and obtain the whole ranking list by sorting these scores in a descending order. This paradigm can greatly improve the inference efficiency as the documents can be scored in parallel. In particular, DALETOR (Yan et al., 2021) and MO4SRD (Yu, 2022) propose a differentiable proxy for the non-differentiable diversity metric $\alpha$-NDCG (Clarke et al., 2008), which enables a gradient-based optimization for the diversity metric. Other methods (Jiang et al., 2017; Qin et al., 2020; Su et al., 2021, 2022; Deng et al., 2024) optimize the diversity of ranking list based on list-pairwise loss fuction (Xia et al., 2008). However, although both the differentiable approximation of the optimization objective and the list-pairwise loss function are good approximations to the true optimization objective, such as alpha-NDCG, they are still not equal to the true optimization objective. Therefore, the solution may still be biased and sub-optimal.
@@
To summarize, although the problem of search result diversification has been extensively investigated, existing methods still have some limitations in terms of the effectiveness and efficiency. From the perspective of effectiveness, the Greedy Selection Approaches (e.g. (Carbonell and Goldstein, 1998; Santos et al., 2010)) and the Simultaneous Scoring and Ranking Approaches (e.g (Yan et al., 2021; Yu, 2022; Jiang et al., 2017; Su et al., 2022)) may be sub-optimal in optimizing the end diversity metric. From the perspective of efficiency, those Single-Agent RL Approaches face a challenge of exploring a huge space of all possible rankings, which inevitably leads to sub-optimal solution (Xia et al., 2017) and high time-complexity in both training and inference (Feng et al., 2018).
@@
In this paper, in order to alleviate the problems of existing methods mentioned above, we propose to formalize the diverse ranking as a multi-agent cooperation process. Since Multi-Agent reinforcement learning (MARL) algorithm is used to optimize the DIVersity ranking model, we call this new method MA4DIV. Specifically, by treating each document as an independent agent within a cooperative multi-agent setting, we simulate a fully cooperative multi-agent task. In this setup, each agent (document) makes action selections based on observations that include the features of both the query and the documents, aiming to maximize a shared reward function that is directly related to the diversity evaluation metrics of search results. Furthermore, MA4DIV employs value decomposition (Rashid et al., 2018) to optimize global diversity directly by the structures of mixing network and hypernetworks during training.
@@
  • Compared to ”greedy selection” paradigm, MA4DIV simultaneously predicts the ranking scores of all documents, which can improve the efficiency of ranking process.
  • The ranking scores of all documents and a ranked documents list can be obtained in one time step. Therefore, MA4DIV has higher exploration efficiency than Single-Agent Reinforcement Learning algorithms, such as MDP-DIV, which must go through an entire episode to explore a diversity ranking list.
  • The training of the MA4DIV model does not require any approximations of different diversity metrics, as it can directly optimize the diversity metrics as rewards during multi-agent reinforcement learning process.
In order to demonstrate the effectiveness and efficiency of MA4DIV, we conducted experiments on TREC benchmark datasets and a new industrial dataset. The experimental results on TREC datasets showed that, MA4DIV achieved the state-of-the-art performance on some evaluation metrics, and training time was significantly shorter than baselines. Considering the small number of queries in TREC dataset (only 198 valid queries), we built a larger diversity dataset based on the real search engine data, which is called DU-DIV. On the DU-DIV dataset, MA4DIV achieved the state-of-the-art performance on all evaluation metrics, and demonstrated a substantial improvement in exploration efficiency and training efficiency.
2. Related Works
@@ 2.1. Search Result Diversification
Carbonell and Goldstein (1998) introduced the maximal marginal relevance criterion, which uses a linear combination of query-document relevance and document novelty to determine the document to be selected. An extension of this concept, the probabilistic latent MMR model (Guo and Sanner, 2010) was proposed by Guo and Sanner. Approaches such as xQuAD (Santos et al., 2010) directly model different aspects of a query, estimating utility based on the relevance of retrieved documents to identified sub-queries or aspects. Hu et al. (2015) proposed a utility function that leverages hierarchical intents of queries, aiming to maximize diversity in the hierarchical structure. Other researchers, like He et al. (2012), have proposed combining implicit and explicit topic representations to create better diverse rankings, while Gollapudi et al. (Gollapudi and Sharma, 2009) proposed an axiomatic approach to result diversification.
@@
Machine learning techniques have also been applied to construct diverse ranking models, often adhering to the sequential document selection or greedy sequential decision-making framework. Several researchers have defined utility as a linear combination of handcrafted relevance and novelty features (Zhu et al., 2014; Xia et al., 2015; Xu et al., 2017). In (Xia et al., 2016), novelty can be modeled using deep learning models such as neural tensor networks. Radlinski et al. (Radlinski et al., 2008) propose to learn a diverse ranking of documents directly based on users’ clicking behavior.
@@
Reinforcement learning algorithms also are proposed to improve the effectiveness of ”greedy selection” paradigm in search result diversification. MDP-DIV (Xia et al., 2017) consider the diverse ranking process as a Markov Decision Process (MDP). In MDP-DIV, agent select one document in one time step, so a ranked list can be obtained after an episode. And the episode reward can be any metrics, such as $\alpha$-DCG. While Feng et al. (Feng et al., 2018) think that ”greedy selection” paradigm is easily lead to local optimum, so $\mathrm{M^{2}Div}$ is proposed to introduce Monte Carlo Tree Search (MCTS) into MDP ranking process. In this way, $\mathrm{M^{2}Div}$ does alleviate the problem that tend to fall into local optimum by conducting high-quality exploration, but also brings expensive training and inference costs.
@@
There are some methods based on score-and-sort paradigm. DALETOR (Yan et al., 2021) and MO4SRD (Yu, 2022) are proposed to derive differentiable diversification-aware losses which are approximation of different diversity metrics, such as $\alpha$-NDCG. And some other methods (Jiang et al., 2017; Qin et al., 2020; Su et al., 2021, 2022; Deng et al., 2024) use list-pairwise loss function (Xia et al., 2008) to optimize the diversity of the ranking list.
@@ 2.2. Reinforcement Learning for IR
Wei et al. (2017) formulate a ranking process as an MDP and train the ranking model with a policy gradient algorithm of REINFORCE (Williams, 1992). Singh et al. (Singh and Joachims, 2019) propose PG Rank algorithm to enhance the fairness of ranking process with reinforcement learning methods. PPG (Xu et al., 2020) uses pairwise comparisons of two sampled document list with in a same query, which makes an unbiased and low variance policy gradient estimations. Yao et al. (2020) model the interactions between search engine and users with a hierarchical Markov Decision Process, which improves the personalized search performance. RLIRank (Zhou and Agichtein, 2020) implement a learning to rank (LTR) model for each iteration of the dynamic search process. Zou et al. (Zou et al., 2019) formulate the ranking process as a multi-agent Markov Decision Process, where the interactions among documents are considered in the ranking process. However, Zou et al. (Zou et al., 2019) conduct the REINFORCE to optimize the ranking model.
@@ 3. Background
3.1. Co-MARL
In this work, we consider the process of diversity ranking as a fully cooperative multi-agent reinforcement learning (Co-MARL) task, which is formally defined as a tuple $G=\left\langle\mathcal{S},\mathcal{A},r,\mathcal{Z},\mathcal{O},n,\gamma\right\rangle$. $s\in\mathcal{S}$ is the state of the environment. Each agent $i\in\mathcal{G}\equiv\{1,\dots,n\}$ chooses an action $a_{i}\in A$ which forms the joint action $\textbf{a}\in\mathcal{A}\equiv A^{n}$. The reward function which is modeled as $r(s,\textbf{a}):\mathcal{S}\times\mathcal{A}$ is shared by all agents and the discount factor is $\gamma\in[0,1)$. In our diversified search task, it follows fully observable settings, where agents have access to the state. Instead, it samples observations $z\in\mathcal{Z}$ according to observation function $\mathcal{O}(s,i):\mathcal{S}\times\mathcal{A}\to\mathcal{Z}$. In our algorithm, the joint policy $\pi$ is based on action-value function $Q^{\pi}_{tot}(s_{t},\textbf{a}_{t})=\mathbb{E}_{s_{t+1}:\infty,\textbf{a}_{t+1}:\infty}[\sum_{k=0}^{\infty}\gamma^{k}r_{t+k}|s_{t},\textbf{a}_{t}]$. The final goal is to get the optimal action-value function $Q^{*}$.
3.2. General Format of Test Set for Diversified Search